# Can Large Multimodal Models Understand Agricultural Scenes? Benchmarking with AgroMind

### v1

@@ 5.1 Evaluation setups
Evaluated Models. We evaluate a total of 21 LMMs using the AgroMind benchmark, including 3 close-sourced and 18 open-sourced models. The close-sourced models including GPT-4o [11], Gemini 1.5 Flash [12], and Claude 3.5 Sonnet [54] are accessed via their official APIs. The open-sourced models fall into two categories: single-image models, including TinyLLaVA [55], the LLaVA-NeXT series [32], Xcomposer2-4KHD [56], InstructBLIP-Vicuna-7B [57], and Idefics-2-8B [58]; and multi-image models, including the DeepSeek-VL2 series [59], InternVL2 series [30], LLaVA-NeXT-Interleave-7B [31], the Mantis series [60], and Idefics-2-8B [58] (see Appendix C.2 for details). All close-sourced models run on machines equipped with NVIDIA A800 GPUs.
@@
GPT-4o [11] 25.71 13.50 20.50 81.62 49.70 45.42 36.81 33.34 45.75 34.00 30.05 58.76 27.39 42.98
Gemini-1.5-Flash [12] 25.00 17.50 28.75 75.42 48.05 65.17 34.64 29.07 39.11 33.06 23.92 60.82 28.06 41.21
Claude-3.5-Sonnet [54] 41.12 17.00 29.00 50.63 45.35 29.50 29.75 22.55 34.64 26.54 14.91 40.09 28.77 33.33
DeepSeek-VL2-tiny [59] 17.86 12.00 16.25 57.88 25.75 45.92 27.88 27.58 12.55 14.02 10.75 37.00 26.98 27.51
DeepSeek-VL2-small [59] 19.88 13.33 21.25 59.91 31.15 37.92 28.36 27.67 23.50 14.35 16.35 31.96 7.20 27.84
TinyLLaVA [55] 35.02 12.00 30.00 63.08 22.20 40.25 26.98 28.60 31.58 7.99 15.90 38.95 10.12 29.88
InternVL2-2B [30] 40.03 15.00 24.75 40.28 27.30 47.50 25.65 25.86 25.73 12.84 28.84 36.88 19.65 28.49
InternVL2-4B [30] 28.59 12.50 28.75 48.18 45.10 13.00 30.54 25.86 39.57 16.02 24.15 41.01 25.52 31.46
XComposer2-4KHD [56] 32.08 11.83 40.00 46.18 37.20 35.75 32.71 27.67 27.11 10.23 38.08 43.53 38.47 32.31
LLaVA-NeXT-7B-Mistral [32] 41.12 12.00 49.25 43.45 42.05 61.00 20.16 28.55 39.93 7.87 11.88 30.93 15.86 31.07
LLaVA-NeXT-7B-Vicuna [32] 7.14 12.00 48.75 47.68 29.75 25.83 18.71 27.76 22.15 9.91 10.22 32.42 20.65 24.61
InstructBLIP-Vicuna-7B [57] 10.84 11.83 14.75 38.09 24.10 16.42 18.47 15.15 15.51 4.32 17.18 22.11 8.08 17.70
LLaVA-NeXT-Interleave-7B [31] 9.42 8.50 25.25 29.75 18.55 3.50 21.00 16.35 19.31 10.84 13.47 23.71 19.40 17.83
Mantis-LLaMA3-SigLIP [60] 20.75 10.83 28.75 63.47 36.70 16.08 28.24 31.32 26.19 12.68 13.55 42.73 36.59 31.60
Mantis-Idefics2 [60] 16.34 11.67 33.75 62.36 46.25 39.92 21.97 27.84 23.65 9.91 14.31 47.65 36.64 31.84
LLaVA-NeXT-8B [32] 45.42 12.17 44.50 51.99 34.75 39.92 25.59 29.63 35.10 8.32 13.10 37.57 25.27 32.07
InternVL2-8B [30] 39.32 11.17 23.50 52.21 41.85 38.25 33.19 28.49 39.50 18.96 16.81 48.34 31.10 34.88
Idefics-2-8b [58] 32.57 13.83 33.50 60.30 36.30 42.58 23.72 27.64 23.59 9.74 14.91 44.10 23.69 30.85
LLaVA-NeXT-13B [32] 38.56 12.00 47.25 55.13 26.70 34.67 20.64 29.95 38.56 5.87 11.13 32.53 31.77 31.43
InternVL2-26B [30] 21.46 13.50 18.25 69.50 43.90 28.33 32.47 28.37 45.13 17.57 12.72 57.62 25.77 35.84
LLaVA-NeXT-34B [32] 48.15 9.33 54.50 57.74 43.00 37.08 22.51 29.10 39.81 13.62 12.72 42.15 27.60 34.93
@@
Evaluation protocols. The AgroMind benchmark comprises five question types: judgment, counting, text choice, image choice and open-ended questions. We follow standard evaluation setups from MMMU [44] and UrBench [17] to parse and assess the answers generated by LMMs. For all question types except open-ended ones, we adopt strict string matching, the answer can be marked correct only if it matches the ground truth exactly. For open-ended questions, correctness is determined based on the semantic similarity between the model-generated answers and the reference answers (see Appendix C.4 for details). Additionally, to ensure that models are evaluated solely based on their own visual understanding rather than leveraging multi-image input advantages, all multi-image questions are converted into single-image input by concatenating all related images into one.
@@ 5.2 Main results
Overall challenge presented in AgroMind. As indicated in Table 2, AgroMind poses significant challenges to current SOTA LMMs. Although some models, such as GPT-4o, have surpassed human performance, 16 out of the 21 evaluated models still underperform compared to humans. Both close-sourced and open-sourced models fall short of consistently outperforming humans. The challenging nature of our benchmark indicates that current LMMs’ strong performance on the standard tasks are not generalized to the multi-modal and multi-scene agricultural scenarios.
@@ 6 Conclusion
In this work, we present AgroMind, a comprehensive benchmark tailored for evaluating LMMs in the context of agricultural remote sensing. The benchmark is constructed from real-world data sources and comprises 25,026 question-answer pairs across 13 task types and 4 progressive cognitive levels: spatial perception, object understanding, scene understanding, and scene reasoning. All tasks are designed to reflect practical agricultural needs, enabling a systematic assessment of LMMs’ ability to interpret, infer, and reason over complex agricultural scenes. We conduct thorough evaluations of 21 representative LMMs, revealing notable performance gaps between current models and expert-level understanding in agriculture-related scenarios. AgroMind not only highlights the limitations of existing models but also provides a new foundation for advancing domain-specific multimodal intelligence. We hope this work inspires future research toward developing more capable, generalizable, and agriculture-aware large models.
@@ B.1 Data collection
Data scene diversity. We carefully select nine datasets—including eight public datasets and one proprietary farmland plot dataset—that cover multiple core scenarios and multimodal data characteristics in agricultural remote sensing. The OAM-TCD Dataset[48] and OilPalmUAV Dataset[49] provide high-resolution global canopy maps and oil-palm tree health assessments, respectively, allowing models to understand vegetation conditions at both macro and micro scales. In addition, the OAM-TCD Dataset[48] encompasses ecologically diverse regions from both the Northern and Southern Hemispheres (e.g., tropical rainforests and temperate woodlands), facilitating precise quantification of forest coverage and area for large-scale forestry monitoring. Meanwhile, the Agriculture Vision Challenge[50] and PhenoBench Dataset[51], by providing pixel-level annotations, emphasize in-field anomaly detection and detailed crop condition analysis, enhancing model capability in handling complex scene details. To support these tasks, the Agriculture Vision Challenge[50] combines RGB and near-infrared channels (512×512 pixels) with six binary masks (e.g., cloud shadow, double plant) to construct a multimodal anomaly diagnosis task. The PhenoBench Dataset[51] utilizes visibility maps (1024×1024 pixels) to quantify plant occlusion levels and frames tasks related to weed coexistence and plant integrity assessment. Furthermore, in the domain of fine-grained cropland management, our 7-Province Ground dataset integrates farmland vector data from multiple provinces with multispectral imagery, enabling precise field boundary delineation and dynamic coverage rate calculation. This supports tasks such as cropland area estimation and boundary recognition, significantly enhancing the practicality of models in agricultural planning and resource optimization.
@@
Multi-sensor heterogeneous data integration. By integrating data from satellite, aerial, and ground-level sensors, this benchmark constructs a relatively comprehensive multi-source heterogeneous dataset for agricultural remote sensing. At the satellite level, the 7-Province Ground Dataset provides multispectral imagery and farmland vector data, supporting field boundary segmentation, area estimation, and coverage rate computation. At the aerial level, the OilPalmUAV[49] Dataset, with high-resolution canopy imagery, serves as the core for low-altitude plant counting and growth status monitoring. Ground-based sensor data further enrich the spectral diversity: PhenoBench Dataset[51] provides hyperspectral data for analyzing plant physiological conditions, while the 2018 AI Challenge Dataset offers disease-annotated leaf images (with 26 labeled disease types), which are paired with severity labels to generate plant disease reasoning QA tasks. To tackle cross-modal fusion challenges, the IP102 Dataset [52], which includes 102 insect categories with VOC-style annotations, is reformulated into a distractor-choice question format to test models’ zero-shot pest identification ability. Through the coordinated use of multiple spatial resolutions, multiple spectral sources (RGB, multispectral, NIR), and diverse annotation formats (pixel-level masks, instance bounding boxes, vector polygons), this dataset enables a systematic evaluation of agricultural LMMs in multimodal data fusion scenarios.
Multi-dimensional evaluation framework. Our benchmark employs a three-dimensional Spatial–Object–Scene framework to establish a hierarchical evaluation system for core agricultural remote-sensing tasks, systematically assessing large-scale multimodal models’ (LMMs) capabilities in localization, comprehension, and reasoning within agricultural contexts. In the Spatial dimension, both global canopy distribution from OAM-TCD Dataset[48] and cropland vector data from the 7-Provinces Ground Dataset support regional area computation and coverage-rate statistics; the latter’s plot-boundary delineation tasks further challenge LMMs’ multi-scale geospatial parsing abilities. In the Object dimension, by leveraging instance-level annotations, the benchmark captures the characteristics of most individual agricultural targets and evaluates models’ precision in detection, classification, and counting within agricultural scenarios. Specifically, the Maize Tassel Identification employs YOLO annotations to construct 3×3 grid–based yield-prediction tasks, validating models’ dense-object localization and associative inference capabilities. Pest-classification tasks in the IP102 Dataset[52] and 2018 AI Challenge Datasets, together with plant-occlusion analysis tasks that PhenoBench Dataset[51] progressively generates from pixel-level visibility annotations, encompass an object-level evaluation spanning basic attribute discrimination to higher-order relational reasoning. In the Scene dimension, we probe models’ capacity for multi-level semantic decomposition, cross-element relational inference, and dynamic decision support in complex agricultural environments. Tasks include multi-scene coverage-rate comparison and distribution inference using OAM-TCD Dataest[48] canopy data, as well as plant-occlusion analysis based on PhenoBench’s pixel-level maps and visibility metrics—requirements that ensure robust understanding and reasoning in interactive scenarios, thereby underpinning reliable decision support for precision agriculture.
@@
Figure 8: Examples of image datasets, with the bar chart showing the quantity of data in different categories.
@@
ACFR Orchard Fruit Dataset [53]     2016     CSV-format annotations for almond, apple, mango; Fruit radius metadata; Quadrant-based distribution analysis     308×202, 500×500, etc.     https://data.acfr.usyd.edu.au/ag/treecrops/2016-multifruit/
@@
IP102 Dataset [52]     2019     102 insect pest categories; VOC-format annotations; Multi-choice QA templates for pest-crop association     Variable     https://github.com/xpwu95/IP102
Agriculture Vision Challenge [50]     2020     Six anomaly types (e.g., Weed Cluster); Multi-spectral data (RGB + NIR); Binary masks for anomaly detection     512×512     https://github.com/SHI-Labs/Agriculture-Vision
OilPalmUAV Dataset [49]     2021     UAV-captured oil palm trees; Instance-level bbox annotations for 5 growth states (Healthy, Dead, Mismanaged, etc.)     1024×1024     https://github.com/rs-dl/MOPAD
PhenoBench Dataset [51]     2023     Hierarchical annotations: semantic segmentation (crop/weed), instance segmentation (plants/leaves), pixel-level visibility maps     1024×1024     https://github.com/PRBonn/phenobench
OAM-TCD Dataset [48]     2024     Global tree cover maps; Semantic segmentation masks; MS-COCO annotations (bbox, polygons); Geospatial metadata (EPSG:3395)     2048×2048     https://zenodo.org/records/11617167
@@
7-Provinces Ground Dataset     2025     Multi-spectral GeoTIFF images; .shp parcel polygons; Reprojected to EPSG:3857; Coverage analysis for farmland     Variable     Internal Collection

### v2

@@ 5.1 Evaluation setups
Evaluated Models. We evaluate a total of 24 LMMs using the AgroMind benchmark, including 4 close-sourced and 20 open-sourced models. The close-sourced models including GPT-4o openai2024gpt4o , Gemini 1.5 Flash team2023gemini , Gemini 1.5 Pro team2023gemini , and Claude 3.5 Sonnet anthropic2024claude3 are accessed via their official APIs. The open-sourced models fall into two categories: single-image models, including TinyLLaVA TinyLLaVA , the LLaVA-NeXT series LLaVA-NeXT , Xcomposer2-4KHD XComposer2 , InstructBLIP-Vicuna-7B InstrucBLIP , Idefics-2-8B Idefics2 , GeoChat kuckreja2023geochat , and GeoLLaVA-8K wang2025geollava ; and multi-image models, including the DeepSeek-VL2 series deepseekvl2 , InternVL2 series InternVL2 , LLaVA-NeXT-Interleave-7B LLaVA-NeXT-Interleave , the Mantis series Mantis , and Idefics-2-8B Idefics2 (see Appendix C.2 for details). All close-sourced models run on machines equipped with NVIDIA A800 GPUs.
@@
GPT-4o openai2024gpt4o 25.71 13.50 20.50 81.62 49.70 45.42 36.81 33.34 45.75 34.00 30.05 58.76 27.39 42.98
Gemini-1.5-Flash team2023gemini 25.00 17.50 28.75 75.42 48.05 65.17 34.64 29.07 39.11 33.06 23.92 60.82 28.06 41.21
Gemini-1.5-Pro team2023gemini 31.69 28.33 30.00 73.95 53.50 30.00 35.98 27.27 40.18 17.55 19.08 68.60 28.33 38.99
Claude-3.5-Sonnet anthropic2024claude3 41.12 17.00 29.00 50.63 45.35 29.50 29.75 22.55 34.64 26.54 14.91 40.09 28.77 33.33
DeepSeek-VL2-tiny deepseekvl2 17.86 12.00 16.25 57.88 25.75 45.92 27.88 27.58 12.55 14.02 10.75 37.00 26.98 27.51
DeepSeek-VL2-small deepseekvl2 19.88 13.33 21.25 59.91 31.15 37.92 28.36 27.67 23.50 14.35 16.35 31.96 7.20 27.84
TinyLLaVA TinyLLaVA 35.02 12.00 30.00 63.08 22.20 40.25 26.98 28.60 31.58 7.99 15.90 38.95 10.12 29.88
InternVL2-2B InternVL2 40.03 15.00 24.75 40.28 27.30 47.50 25.65 25.86 25.73 12.84 28.84 36.88 19.65 28.49
InternVL2-4B InternVL2 28.59 12.50 28.75 48.18 45.10 13.00 30.54 25.86 39.57 16.02 24.15 41.01 25.52 31.46
XComposer2-4KHD XComposer2 32.08 11.83 40.00 46.18 37.20 35.75 32.71 27.67 27.11 10.23 38.08 43.53 38.47 32.31
LLaVA-NeXT-7B-Mistral LLaVA-NeXT 41.12 12.00 49.25 43.45 42.05 61.00 20.16 28.55 39.93 7.87 11.88 30.93 15.86 31.07
LLaVA-NeXT-7B-Vicuna LLaVA-NeXT 7.14 12.00 48.75 47.68 29.75 25.83 18.71 27.76 22.15 9.91 10.22 32.42 20.65 24.61
InstructBLIP-Vicuna-7B InstrucBLIP 10.84 11.83 14.75 38.09 24.10 16.42 18.47 15.15 15.51 4.32 17.18 22.11 8.08 17.70
LLaVA-NeXT-Interleave-7B LLaVA-NeXT-Interleave 9.42 8.50 25.25 29.75 18.55 3.50 21.00 16.35 19.31 10.84 13.47 23.71 19.40 17.83
Mantis-LLaMA3-SigLIP Mantis 20.75 10.83 28.75 63.47 36.70 16.08 28.24 31.32 26.19 12.68 13.55 42.73 36.59 31.60
Mantis-Idefics2 Mantis 16.34 11.67 33.75 62.36 46.25 39.92 21.97 27.84 23.65 9.91 14.31 47.65 36.64 31.84
LLaVA-NeXT-8B LLaVA-NeXT 45.42 12.17 44.50 51.99 34.75 39.92 25.59 29.63 35.10 8.32 13.10 37.57 25.27 32.07
InternVL2-8B InternVL2 39.32 11.17 23.50 52.21 41.85 38.25 33.19 28.49 39.50 18.96 16.81 48.34 31.10 34.88
Idefics-2-8b Idefics2 32.57 13.83 33.50 60.30 36.30 42.58 23.72 27.64 23.59 9.74 14.91 44.10 23.69 30.85
LLaVA-NeXT-13B LLaVA-NeXT 38.56 12.00 47.25 55.13 26.70 34.67 20.64 29.95 38.56 5.87 11.13 32.53 31.77 31.43
InternVL2-26B InternVL2 21.46 13.50 18.25 69.50 43.90 28.33 32.47 28.37 45.13 17.57 12.72 57.62 25.77 35.84
LLaVA-NeXT-34B LLaVA-NeXT 48.15 9.33 54.50 57.74 43.00 37.08 22.51 29.10 39.81 13.62 12.72 42.15 27.60 34.93
GeoChat kuckreja2023geochat 32.79 8.33 30.00 44.26 27.00 38.33 20.73 24.05 16.56 8.57 18.32 41.86 27.50 26.15
GeoLLaVA-8K wang2025geollava 9.29 11.67 12.50 27.45 21.00 10.83 18.29 8.21 19.63 7.76 9.92 29.07 7.92 15.24
@@
Evaluation protocols. The AgroMind benchmark comprises five question types: judgment, counting, text choice, image choice and open-ended questions. We follow standard evaluation setups from MMMU yue2024mmmu and UrBench zhou2025urbench to parse and assess the answers generated by LMMs. For all question types except open-ended ones, we adopt strict string matching, the answer can be marked correct only if it matches the ground truth exactly. For open-ended questions, correctness is determined based on the semantic similarity between the model-generated answers and the reference answers (see Appendix C.4 for details). Additionally, to ensure that models are evaluated solely based on their own visual understanding rather than leveraging multi-image input advantages, all multi-image questions are converted into single-image input by concatenating all related images into one.
@@ 5.2 Main results
Overall challenge presented in AgroMind. As indicated in Table 2, AgroMind poses significant challenges to current SOTA LMMs. Although some models, such as GPT-4o, have surpassed human performance, 18 out of the 24 evaluated models still underperform compared to humans. Both close-sourced and open-sourced models fall short of consistently outperforming humans. The challenging nature of our benchmark indicates that current LMMs’ strong performance on the standard tasks are not generalized to the multi-modal and multi-scene agricultural scenarios. Additionally, domain-specific models like GeoChat kuckreja2023geochat and GeoLLaVA-8K wang2025geollava , while effective in general remote sensing tasks, demonstrate limitations and reduced generalization in complex agricultural scenarios.
@@ 6 Conclusion
In this work, we present AgroMind, a comprehensive benchmark tailored for evaluating LMMs in the context of agricultural remote sensing. The benchmark is constructed from real-world data sources and comprises 27,247 question-answer pairs across 13 task types and 4 progressive cognitive levels: spatial perception, object understanding, scene understanding, and scene reasoning. All tasks are designed to reflect practical agricultural needs, enabling a systematic assessment of LMMs’ ability to interpret, infer, and reason over complex agricultural scenes. We conduct thorough evaluations of 24 representative LMMs, revealing notable performance gaps between current models and expert-level understanding in agriculture-related scenarios. AgroMind not only highlights the limitations of existing models but also provides a new foundation for advancing domain-specific multimodal intelligence. We hope this work inspires future research toward developing more capable, generalizable, and agriculture-aware large models.
@@ B.1 Data collection
Figure 8: Geographical coverage map of datasets.
Geospatial distribution. To address spatial coverage concerns, we conducted statistical analysis of global data distribution (see Fig. 8). AgroMind intentionally spans 80 geographic regions including China, the United States, European Union nations, Brazil, Southeast Asia, and parts of Africa, reflecting agricultural diversity across crops, climates, and land-use patterns, achieving an extensive spatial coverage to date for agricultural remote sensing benchmarks.This global representation captures critical agricultural diversity across: (i) Climate regimes – from tropical monsoons (Indonesia) to Mediterranean (Italy) and continental (Ukraine); (ii) Land-use intensity – ranging from industrial monocultures (U.S. Corn Belt) to polycultures (Tanzanian farmlands); (iii) Crop phenology – covering growth cycles of many staple crops (maize, rice, oil palm, etc.). The rich geographic diversity of the dataset allows robust cross-regional generalization validation and provides foundational resources for global agricultural sustainability research.
@@
Data scene diversity. We carefully select nine datasets—including eight public datasets and one proprietary farmland plot dataset—that cover multiple core scenarios and multimodal data characteristics in agricultural remote sensing. The OAM-TCD Datasetveitchmichaelis2024oamtcdgloballydiversedataset and OilPalmUAV Datasetzheng2021growing provide high-resolution global canopy maps and oil-palm tree health assessments, respectively, allowing models to understand vegetation conditions at both macro and micro scales. In addition, the OAM-TCD Datasetveitchmichaelis2024oamtcdgloballydiversedataset encompasses ecologically diverse regions from both the Northern and Southern Hemispheres (e.g., tropical rainforests and temperate woodlands), facilitating precise quantification of forest coverage and area for large-scale forestry monitoring. Meanwhile, the Agriculture Vision ChallengeChiu_2020_CVPR and PhenoBench Datasetweyler2024pami , by providing pixel-level annotations, emphasize in-field anomaly detection and detailed crop condition analysis, enhancing model capability in handling complex scene details. To support these tasks, the Agriculture Vision ChallengeChiu_2020_CVPR combines RGB and near-infrared channels (512×512 pixels) with six binary masks (e.g., cloud shadow, double plant) to construct a multimodal anomaly diagnosis task. The PhenoBench Datasetweyler2024pami utilizes visibility maps (1024×1024 pixels) to quantify plant occlusion levels and frames tasks related to weed coexistence and plant integrity assessment. Furthermore, in the domain of fine-grained cropland management, our 7-Province Ground dataset integrates farmland vector data from multiple provinces with multispectral imagery, enabling precise field boundary delineation and dynamic coverage rate calculation. This supports tasks such as cropland area estimation and boundary recognition, significantly enhancing the practicality of models in agricultural planning and resource optimization.
@@
Multi-sensor heterogeneous data integration. By integrating data from satellite, aerial, and ground-level sensors, this benchmark constructs a relatively comprehensive multi-source heterogeneous dataset for agricultural remote sensing. At the satellite level, the 7-Province Ground Dataset provides multispectral imagery and farmland vector data, supporting field boundary segmentation, area estimation, and coverage rate computation. At the aerial level, the OilPalmUAVzheng2021growing Dataset, with high-resolution canopy imagery, serves as the core for low-altitude plant counting and growth status monitoring. Ground-based sensor data further enrich the spectral diversity: PhenoBench Datasetweyler2024pami provides hyperspectral data for analyzing plant physiological conditions, while the 2018 AI Challenge Dataset offers disease-annotated leaf images (with 26 labeled disease types), which are paired with severity labels to generate plant disease reasoning QA tasks. To tackle cross-modal fusion challenges, the IP102 Dataset Wu2019Insect , which includes 102 insect categories with VOC-style annotations, is reformulated into a distractor-choice question format to test models’ zero-shot pest identification ability. Through the coordinated use of multiple spatial resolutions, multiple spectral sources (RGB, multispectral, NIR), and diverse annotation formats (pixel-level masks, instance bounding boxes, vector polygons), this dataset enables a systematic evaluation of agricultural LMMs in multimodal data fusion scenarios.
@@
Figure 9: Examples of image datasets, with the bar chart showing the quantity of data in nine detailed subclasses. These subclasses refine the five macro-categories shown in Fig. 3 to microscopically validate task diversity.
Multi-dimensional evaluation framework. Our benchmark employs a three-dimensional Spatial–Object–Scene framework to establish a hierarchical evaluation system for core agricultural remote-sensing tasks, systematically assessing large-scale multimodal models’ (LMMs) capabilities in localization, comprehension, and reasoning within agricultural contexts. In the Spatial dimension, both global canopy distribution from OAM-TCD Datasetveitchmichaelis2024oamtcdgloballydiversedataset and cropland vector data from the 7-Provinces Ground Dataset support regional area computation and coverage-rate statistics; the latter’s plot-boundary delineation tasks further challenge LMMs’ multi-scale geospatial parsing abilities. In the Object dimension, by leveraging instance-level annotations, the benchmark captures the characteristics of most individual agricultural targets and evaluates models’ precision in detection, classification, and counting within agricultural scenarios. Specifically, the Maize Tassel Identification employs YOLO annotations to construct 3×3 grid–based yield-prediction tasks, validating models’ dense-object localization and associative inference capabilities. Pest-classification tasks in the IP102 DatasetWu2019Insect and 2018 AI Challenge Datasets, together with plant-occlusion analysis tasks that PhenoBench Datasetweyler2024pami progressively generates from pixel-level visibility annotations, encompass an object-level evaluation spanning basic attribute discrimination to higher-order relational reasoning. In the Scene dimension, we probe models’ capacity for multi-level semantic decomposition, cross-element relational inference, and dynamic decision support in complex agricultural environments. Tasks include multi-scene coverage-rate comparison and distribution inference using OAM-TCD Dataestveitchmichaelis2024oamtcdgloballydiversedataset canopy data, as well as plant-occlusion analysis based on PhenoBench’s pixel-level maps and visibility metrics—requirements that ensure robust understanding and reasoning in interactive scenarios, thereby underpinning reliable decision support for precision agriculture.
Temporal and phenological coverage. AgroMind establishes a comprehensive temporal axis essential for modeling crop phenodynamics, systematically capturing growth transitions, anomaly evolution, and seasonal patterns through multi-source heterogeneous data. Specifically: (i) Growth-stage representation leverages the OilPalmUAVzheng2021growing Dataset to construct QA pairs tracking developmental phases from growth ,maturity to aging, while competition management tasks at critical phenophases utilize PhenoBench’s visibility metricsweyler2024pami to quantify weed-crop interactions; (ii) Anomaly temporal sensitivity encodes time-critical states through spectral signatures—waterlogging detection via near-infrared absorption characteristics in Agriculture Vision ChallengeChiu_2020_CVPR —enabling cross-temporal implicit comparison; (iii) Multi-seasonal imagery combines the cultivated land status of multiple provinces in China at different time series (7 provincial ground datasets) and the Global Ecoregion Tree Monitoring Imagery (OAM-TCDveitchmichaelis2024oamtcdgloballydiversedataset ), covering the crop cycle to support phenological transition modeling. This temporal richness enables evaluation of longitudinal reasoning capabilities, particularly for yield prediction and climate adaptation analysis, establishing a foundation for multi-temporal agricultural assessment.
@@
ACFR Orchard Fruit Dataset bargoti2016deep     2016     CSV-format annotations for almond, apple, mango; Fruit radius metadata; Quadrant-based distribution analysis     308×202, 500×500, etc.     https://data.acfr.usyd.edu.au/ag/treecrops/2016-multifruit/
@@
IP102 Dataset Wu2019Insect     2019     102 insect pest categories; VOC-format annotations; Multi-choice QA templates for pest-crop association     Variable     https://github.com/xpwu95/IP102
Agriculture Vision Challenge Chiu_2020_CVPR     2020     Six anomaly types (e.g., Weed Cluster); Multi-spectral data (RGB + NIR); Binary masks for anomaly detection     512×512     https://github.com/SHI-Labs/Agriculture-Vision
OilPalmUAV Dataset zheng2021growing     2021     UAV-captured oil palm trees; Instance-level bbox annotations for 5 growth states (Healthy, Dead, Mismanaged, etc.)     1024×1024     https://github.com/rs-dl/MOPAD
PhenoBench Dataset weyler2024pami     2023     Hierarchical annotations: semantic segmentation (crop/weed), instance segmentation (plants/leaves), pixel-level visibility maps     1024×1024     https://github.com/PRBonn/phenobench
OAM-TCD Dataset veitchmichaelis2024oamtcdgloballydiversedataset     2024     Global tree cover maps; Semantic segmentation masks; MS-COCO annotations (bbox, polygons); Geospatial metadata (EPSG:3395)     2048×2048     https://zenodo.org/records/11617167
@@
7-Provinces Ground Dataset     2025     Multi-spectral GeoTIFF images; .shp parcel polygons; Reprojected to EPSG:3857; Coverage analysis for farmland (After further supplementation, the distribution of the dataset covers not only seven provinces in China, but also some regions in Southeast Asia and Europe.)     Variable     Internal Collection